OpenStudy (anonymous):

may any one help me in this exercise if p(t) = 9.865(1.025)^t find p'(t)

OpenStudy (mathmate):

hint: \( \frac{d\ a^t}{dt}\ = \ a^t \ ln(t) \) set a=1.025 note that 9.865 is just a multiplicative constant.

OpenStudy (anonymous):

\[\frac{d\ a^t}{dt}\ = \ a^t \ \ln(t)*t'\]